A model of a helicopter rotor has four blades, each 3.40 m long from the central shaft to the blade tip. The model is rotated in a wind tunnel at 550 rev/min. What is the radial acceleration (in m/s2) of the blade tip?


1
Expert's answer
2022-06-06T08:53:35-0400

The radial acceleration

"a=\omega^2R=(2\pi n)^2R\\
=(2*3.14*550/60)^2*3.40=1.23*10^4\:\rm m/s^2"